OpenStudy (anonymous):

Find the Surface Area of a cone with a diameter of 10 cm and a height of 12 cm. Use pi = 3.14.

OpenStudy (anonymous):

\[SA=\pi*rs+\pi*r^2\] The radius would be 5 cm and so the slant height would be 13 cm Thus the SA=90*pi or 282.74

OpenStudy (anonymous):

Total surface area of the cone: \[SA=\pi*r ^{2}+\pi*r*\sqrt({r}^{2}+h ^{2})\] plug-in your numbers :)

OpenStudy (amistre64):

i get: \[Sa=pi\ r\sqrt{r^2+h^2}\]

OpenStudy (amistre64):

but thats not accounting for the area of the base is it...

OpenStudy (anonymous):

ok so 3.14*5 * square root of 25+144?

OpenStudy (amistre64):

yeah; which is 65pi ; but methinks thats just the lateral area; if we include the base add another 25pi?

OpenStudy (anonymous):

15.7*13 so the final answer is 204.1?

OpenStudy (anonymous):

if im using 3.14 as pi

OpenStudy (amistre64):

90*3.14 = 282.6 ?

OpenStudy (anonymous):

wait where did u get 90?

OpenStudy (amistre64):

i tend to confuse lateral area and surface area; cause im old :)

OpenStudy (amistre64):

65 pi is the lateral area; i think surface area adds the base area which is another 25 pi

OpenStudy (anonymous):

To get the lateral area you multiply the slant area by the radius by pi. However, you also have to add the base area which is pi* radius squared.

OpenStudy (anonymous):

So amistre is right, it is 65pi+25pi=90pi

OpenStudy (anonymous):

so did i do that right with the formula that was given above?

OpenStudy (anonymous):

im looking for surface area though

OpenStudy (anonymous):

Yes the surface area is the lateral area + the base area which is 90 pi

OpenStudy (amistre64):

65*3.14 is 204.1; but thats not the whole surface area ....

OpenStudy (anonymous):

then what is?

OpenStudy (amistre64):

add another 25 pi to it to include the area of the base 90pi altogether

OpenStudy (amistre64):

282.6

OpenStudy (amistre64):

pi r^2 = area of the base pi 5^2 = 25pi

OpenStudy (anonymous):

i am very confused lol can u show me the steps?

OpenStudy (amistre64):

Add the Area of the sides and the Area of the Base to get total surface area

OpenStudy (amistre64):

\[Sa=(3.14(r)*\sqrt{r^2+h^2})\ +\ (\pi \ r^2)\]

OpenStudy (anonymous):

so 3.14*25 + 3.14*10 * sqrt 25+144?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

3171.4*13=41228.2 ?

OpenStudy (amistre64):

you hit a wrong key i think :) 3.14(5^2) + 3.14(5)(13)

OpenStudy (amistre64):

282.6 is what you should get

OpenStudy (anonymous):

oh i got it now lol my bad

OpenStudy (amistre64):